OpenStudy (anonymous):

Two vectors, A = 10i + 15j + 20k and B = 12i – 6j + k are given. What is the magnitude of C = 4A - 5B What is the scalar product AB? What is the angle between A and B? What is the vector product AxB?

OpenStudy (anonymous):

What are you having troubles with?

OpenStudy (anonymous):

see i know how to the magnitude thing but can you help with th scalar product

OpenStudy (anonymous):

The scalar product, also known as the dot product, is the product defined by multiplying the corresponding components of each vector and adding them together.

OpenStudy (anonymous):

So take vectors i and j. i * j is zero as 1*0 + 0*1 + 0*0 = 0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Ok. Let's say we have two vectors \[x = \left(\begin{matrix}1 \\ 2 \\ 3\end{matrix}\right) y = \left(\begin{matrix}0 \\ 1 \\ -1\end{matrix}\right)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Then the scalar product, x*y, is \[x * y = 1*0 + 2*1 + 1*(-1) = \]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

I mean x*y = 1*0 + 2*1 + 3 *(-1) = -1

OpenStudy (anonymous):

ok lets start will the magnitude one \[A= 50i + 25j + 40k\] \[B = 42i -16j + 34k\] \[C = 2A - B\]

OpenStudy (anonymous):

find the magnitude of A and B

OpenStudy (anonymous):

Do you need help with calculating the magnitude?

OpenStudy (anonymous):

can you help with calculating magnitude of A and B and for C i got 58i+66j+46k

OpenStudy (anonymous):

C should be 58i + 36j + 46k

OpenStudy (anonymous):

And the magnitude of a vector is the sqrt of the sum of its components squared. so for A, |A| = sqrt( 50^2 + 25^2 + 40^2 ) = sqrt(4752)

OpenStudy (anonymous):

so the answer would be 68.738

OpenStudy (anonymous):

for C this square root will be our next method

OpenStudy (anonymous):

Nextmethod?

OpenStudy (anonymous):

since we had found C now we will find |C|

OpenStudy (anonymous):

Ok. So we just do what we did with A.

OpenStudy (anonymous):

squart(50^2 + 25^ 2 + 40^2 )

OpenStudy (anonymous):

We wouldn't use the components of A. We use the components of C.